The Debtors List given below was extracted from the books of Sippy Stores on 31 March 2020, before some errors were discovered: The following errors must now be taken into account to correct the debtors list: 1. A credit sale of goods for R640 to H. Hammond was erroneously posted to his account as R460. 2. A credit invoice issued to G. Fritz for goods sold was erroneously recorded in his account as sales returns, R600. 3. Interest charged at 12% p.a. on R. Roman’s account that was overdue for three months was not posted to his account. 1.7 The correct balance of H. Hammond’s account is _________. (3) A R2 040 B R1 220 C R1 580 D 1480


1
Expert's answer
2021-09-20T12:01:46-0400

Solution:

The correct answer is D). R1,480

The initial journal entry was as follows:

Dr. Account receivable A/c                         460

                Cr. Sales A/c                                                        460

Actual sales amount = 640

Understated amount = 640 – 460 = 180

The account for H. Hammond was understated by 180.

Correcting journal entry is as follows:

Dr. Account Receivable A/c                       180

               Cr. Sales A/c                                                        180

(Correcting entry for understatement)

H. Hammond initial account balance = 1,300

Add understatement of 180 to reflect the actual balance

The correct balance of H. Hammond’s account is = 1,300 + 180 = 1,480